OpenStudy (anonymous):

eric can wash the family van is 1hr. stefan can wash it in 45 min. how long will it take them if they work together?

OpenStudy (unklerhaukus):

w=60mins eric w=45 mins stefan w/60=e w/45=s e+s=w/60+w/45=

OpenStudy (unklerhaukus):

simplify and solve for w

OpenStudy (unklerhaukus):

how many minutes?